Two long, straight, parallel wires 24 cm apart carry currents of I1 = 30A and
I2 = 20A in the same direction. Determine the magnetic field in the plane of the two wires at a point halfway between the wires.
1
Expert's answer
2019-08-28T09:18:38-0400

The magnitude of magnetic field is given by formula

"B=\frac{µ_{0} ×I_{1}}{πd} - \frac{µ_{0} × I_{2}}{πd} (1)"

where µ0 is magnetic constant in vacuum

In our case, we have I1 = 30A, I2 = 20A, d=24 cm

Using (1) we got

B=1.67×10-5 T.


Answer

1.67×10-5 T